A heat engine does 23 kJ of work during 1800 s. Find the power supplied by the engine.


1
Expert's answer
2021-05-18T11:10:15-0400
"P=\dfrac{W}{t}=\dfrac{23\cdot10^3\ J}{1800\ s}=12.8\ W."